OpenStudy (anonymous):

Let f(x) = x – 3 and g(x) = x + 11. Find f(x) • g(x).

OpenStudy (anonymous):

answer choices: x2 – 8x – 33 x2 – 8x + 14 x2 + 8x + 14 x2 + 8x – 33

OpenStudy (anonymous):

f(x) = x - 3 g(x) = x + 11 Find f(x) • g(x) =(x - 3) * (x + 11)

OpenStudy (anonymous):

= x^2 - 3x + 11x - 33 = x^2 + 8x - 33